Construct a 99% confidence interval for the population standard deviation σ if a sample of size 8 has standard deviation 


1
Expert's answer
2021-08-24T17:08:36-0400

We need to construct the 99% confidence interval for the population variance. We have been provided with the following information about the sample variance and sample size:


"s=7.5, s^2=7.5^2=56.25, n=8."

The critical values for "\alpha=0.01" and "df=n-1=7" degrees of freedom are "\chi_L^2=\chi^2_{1-\alpha/2, n-1}=0.9893"  and"\chi_U^2=\chi^2_{\alpha/2, n-1}=20.2777."

The corresponding confidence interval is computed as shown below:


"CI(Variance)=(\dfrac{(n-1)s^2}{\chi^2_{\alpha/2, n-1}}, \dfrac{(n-1)s^2}{\chi^2_{1-\alpha/2, n-1}})"

"=(\dfrac{(8-1)56.25}{20.2777}, \dfrac{(8-1)56.25}{0.9893})"

"\approx(19.4178, 398.0265)"


"CI(Standard \ Deviation)=(\sqrt{19.4178}, \sqrt{398.0265})"

"\approx(4.4066, 19.9506)"